Please type your answer! don't do hand written please

This is the given question by instructor
Answer #1

solution: First calculate the pump work. The specific gravity is 1.03 and the change in pressure is 5.95 MPa, so the change in enthalpy is (1.03)*(5.95)=6.13. Then the value of h2 is h1 + wp =34.5+6.1 = 346.6.

Next, at 50 kPa, the saturated liquid entropy is 1.0910 and that of saturated vapor is 7.5939, and the entropy of the vapor before and after expansion in the turbine is s3 =s4 =5.8892, so the quality x is calculated as follows:

x= (5.8892-1.0910)/(7.5939-1.0910) = 0.738

Since h1 = 340.5 and at 50 kPa the enthalpy of saturated vapor is 2645.9, the value of h4 is calculated as follows:

h4 = h1 + x(hg-h1) = 340.5 - 0.738 (2645.9 - 340.5) =2041.6

The values of h and s for 1 to 4

State h s

1 340.5 Not needed

2 346.6 Not needed

3 2784.3 5.8892

4 2041.6 5.8892

Heat input into the cycle is h3 - h2 =2437.7. Work output is h3 - h4 = 742.7. After deducing pump work, net work output is 736.6. Thermal efficiency is therefore

  \etath = Wnet /Qnet =736.6 /2437.7 = 0.302 =30.2%